LadyRavyRay:

Solve the equation. You may find using algebra tiles helpful in solving this problem. x + 2 = −x − 2 x =

XxAngelGirlxX:

Hi! do you know how to solve this?

LadyRavyRay:

Not really.

XioGonz:

x+2-(-x-2)=0 2x + 4 = 2 • (x + 2) Solve : 2 = 0 Solve : x+2 = 0 Subtract 2 from both sides of the equation x = -2

XxAngelGirlxX:

Xio got it bedore me lol!

LadyRavyRay:

Um...-2? I have no idea

XioGonz:

lol Good job though

XxAngelGirlxX:

yes x= -2

Laylalyssa:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @XxAngelGirlxX Xio got it bedore me lol! \(\color{#0cbb34}{\text{End of Quote}}\) bedore?

elpapitocay:

.

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Laylalyssa \(\color{#0cbb34}{\text{Originally Posted by}}\) @XxAngelGirlxX Xio got it bedore me lol! \(\color{#0cbb34}{\text{End of Quote}}\) bedore? \(\color{#0cbb34}{\text{End of Quote}}\) LMAo

XxAngelGirlxX:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz \(\color{#0cbb34}{\text{Originally Posted by}}\) @Laylalyssa \(\color{#0cbb34}{\text{Originally Posted by}}\) @XxAngelGirlxX Xio got it bedore me lol! \(\color{#0cbb34}{\text{End of Quote}}\) bedore? \(\color{#0cbb34}{\text{End of Quote}}\) LMAo \(\color{#0cbb34}{\text{End of Quote}}\) before* im half asleep